OpenStudy (anonymous):

A family has two cars. The first car has a fuel efficiency of 35 miles per gallon of gas and the second has a fuel efficiency of 20 miles per gallon of gas. During one particular week, the two cars went a combined total of 1575 miles, for a total gas consumption of 60 gallons. How many gallons were consumed by each of the two cars that week?

OpenStudy (dtan5457):

quick note that's not the best way to format your question lol

OpenStudy (michele_laino):

I think if I call with x number of gallon used by the first car and with y number og gallon used by the second car, then from your data I can write this: \[35*x+20*y=1575\] and \[x+y=60\] I simply mathematized your problem text!

OpenStudy (dtan5457):

Going on from @Michele_Laino Now, solve it like a system of equations 35x+20y=1575 x+y=60 ^^^^^ smart choice to isolate one of the variables on the 2nd equation

OpenStudy (dtan5457):

To plug into the first ^

OpenStudy (dtan5457):

Do you know how to do that?

OpenStudy (dtan5457):

@autumnshadow